■ TABLE 62 Compressive Strength (in psi) of 80 Aluminum-Lithium Alloy Specimens 105 97 245 163 207 134 218 199 160 196 221 154 228 131 180 178 157 151 175 201 183 153 174 154 190 76 101 142 149 200 186 174 199 115 193 167 121 120 181 160 194 184 165 145 160 150 181 168 158 208 133 135 172 180 167 176 158 156 229 158 148 150 118 143 141 110 133 123 146 169 158 135 149 163 87 176 237 170

Consider the compressive strength data 6-2. What propotion of thespecimens exhibit compressive strength of at least 1379 KN/m^2?
Answer #1

1379 KN/m2 = 200.007 psi

From the given data of 80 values,

Number of values \geq 200.007 = 9

So,

Proportion at least 1379 KN/m2   = 9/80 = 0.1125

So,

Answer is:

0.1125